# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

### v2

@@ V Conclusion
In this work, we presented FRNet, a Frustum-Range network designed for efficient and effective LiDAR segmentation, which can be seamlessly integrated into range-view approaches. FRNet comprises three key components: the Frustum Feature Encoder Module for extracting per-point features at lower levels, the Frustum-Point Fusion Module for hierarchical updating of per-point features, and the Fusion Head Module for fusing features from different levels to predict more accurate labels. Additionally, We introduced two efficient augmentation methods, FurstumMix and RangeInterpolation, to enhance the generalizability of range-view LiDAR segmentation. Extensive experiments across diverse driving perception datasets verified that FRNet achieves competitive performance while maintaining high efficiency. We hope this work can lay a solid foundation for future works targeted at real-time, in-vehicle 3D perception. Moving forward, we will further enhance the representation learning capability of FRNet and incorporate such a method into other 3D perception tasks, such as 3D object detection and semantic occupancy prediction.
@@ VI-A Scalability and Robustness
TABLE VI: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [28] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE VII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [43]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE VIII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [43]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE IX: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [63] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE X: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [64] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@ VI-B1 SemanticKITTI
Tab. VI shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [28]. We compare FRNet with different LiDAR representations, including point-view, range-view, sparse-voxel-view, and multi-view methods. Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods. Compared with SphereFormer [68] and UniSeg [23], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy. More results about efficiency and accuracy can be found in the main body. It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
@@
Tab. VII and Tab. VIII show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [43], respectively. The results demonstrate the great advantages of Fast-FRNet and FRNet. For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters. Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
@@ VI-B3 ScribbleKITTI
To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [63], including RangeNet++ [30], SalsaNext [57], FIRNet [55], CENet [58], and RangeViT [31]. All the models are trained with their official settings. As shown in Tab. IX, FRNet achieves state-of-the-art performance. However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
@@ VI-B4 SemanticPOSS
We also reimplement some popular sparse-voxel-view methods, including SPVCNN [15], Cylinder3D [16], and MinkNet [17], on SemanticPOSS [64]. All the experiments are conducted based on MMDetection3D framework [72]. Tab. X summarizes the class-wise IoU scores among these methods and some other popular range-view methods. Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods. For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.

### v3

@@ V Conclusion
In this work, we presented FRNet, a Frustum-Range network designed for efficient and effective LiDAR segmentation, which can be seamlessly integrated into range-view approaches. FRNet comprises three key components: the Frustum Feature Encoder Module for extracting per-point features at lower levels, the Frustum-Point Fusion Module for hierarchical updating of per-point features, and the Fusion Head Module for fusing features from different levels to predict more accurate labels. Additionally, we introduced two efficient augmentation methods, FurstumMix and RangeInterpolation, to enhance the generalizability of range-view LiDAR segmentation. Extensive experiments across diverse driving perception datasets verified that FRNet achieves competitive performance while maintaining high efficiency. We hope this work can lay a solid foundation for future works targeted at real-time, in-vehicle 3D perception. Moving forward, we will further enhance the representation learning capability of FRNet and incorporate such a method into other 3D perception tasks, such as 3D object detection and semantic occupancy prediction, and comprehensive vehicle-road coordination perception tasks.
@@ VI-A Scalability and Robustness
TABLE X: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [32] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE XI: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [47]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE XII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [47]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE XIII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [67] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
TABLE XIV: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [68] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@ VI-B1 SemanticKITTI
Tab. X shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [32]. We compare FRNet with different LiDAR representations, including point-view, range-view, sparse-voxel-view, and multi-view methods. Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods. Compared with SphereFormer [72] and UniSeg [24], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy. More results about efficiency and accuracy can be found in the main body. It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
@@
Tab. XI and Tab. XII show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [47], respectively. The results demonstrate the great advantages of Fast-FRNet and FRNet. For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters. Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
@@ VI-B3 ScribbleKITTI
To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [67], including RangeNet++ [34], SalsaNext [64], FIRNet [59], CENet [61], and RangeViT [35]. All the models are trained with their official settings. As shown in Tab. XIII, FRNet achieves state-of-the-art performance. However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
@@ VI-B4 SemanticPOSS
We also reimplement some popular sparse-voxel-view methods, including SPVCNN [15], Cylinder3D [16], and MinkNet [17], on SemanticPOSS [68]. All the experiments are conducted based on MMDetection3D framework [77]. Tab. XIV summarizes the class-wise IoU scores among these methods and some other popular range-view methods. Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods. For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.